Suppose X~Binomial(n,p).Find the pgf of Y=X+1 in terms of GP(x)

1
Expert's answer
2021-12-14T12:57:50-0500

The probability generating function (PGF) of X is GX(s) = E(sX), for all s ∈ R for which the sum converges.

for Binomial Distribution:


"G_X(s)=\sum s^x\begin{pmatrix}
 n \\
 x
\end{pmatrix}p^xq^{n-x}=\sum \begin{pmatrix}
 n \\
 x
\end{pmatrix}(ps)^xq^{n-x}=(ps+q)^n"


"G_Y(s)=\sum \begin{pmatrix}
 n \\
 y
\end{pmatrix}(ps)^yq^{n-y}=\sum \begin{pmatrix}
 n \\
 x+1
\end{pmatrix}(ps)^{x+1}q^{n-(x+1)}=(ps+q)^n=G_X(s)"